### v2

# supplement.pdf

Abstract
Anderson localization physics features three fundamental types of eigenstates: extended, localized, and critical. Confirming the presence of critical states necessitates either advancing the analysis to the thermodynamic limit or identifying a universal mechanism which can rigorously determine these states. Here we report the unambiguous experimental realization of critical states, governed by a rigorous mechanism for exact quantum critical states, and further observe a generalized mechanism that quasiperiodic zeros in hopping couplings protect the critical states. Leveraging a superconducting quantum processor with up to 56 qubits, we implement a programmable mosaic model with tunable couplings and on-site potentials. By measuring time-evolved observables, we identify both delocalized dynamics and incommensurately distributed zeros in the couplings, which are the defining features of the critical states. We map the localized-to-critical phase transition and demonstrate that critical states persist until quasiperiodic zeros are removed by strong long-range couplings, highlighting a novel generalized mechanism discovered in this experiment and shown with rigorous theory. Finally, we resolve the energy-dependent transition between localized and critical states, revealing the presence of anomalous mobility edges.
Introduction. In an ideal crystal, electrons experience a spatially periodic potential and are characterized by Bloch states that possess lattice translational symmetry, extending uniformly throughout the material. Introducing disorder breaks the translational symmetry, impeding the extension of the wave functions, thereby localizing the electronic states; this is known as Anderson localization 1, 2, 3, 4, 5. The states at the extended-localized transition, however, are neither fully extended nor localized, but are in a critical phase. The quantum critical phase is one of the fundamental phases in Anderson localization physics 6, 7, 8, 9, 10, 11, surpassing the complexity of the extended and localized phases 12, 13, 14, 15, 16, 17, and has attracted considerable interest 18. This phase manifests as delocalized matter waves in both position and momentum spaces with local scale invariance as shown in Fig. 1a-c, reflecting the interplay between self-duality and multifractal structures 19, 20, 21, 22, 23. The inclusion of interactions further enriches the physics of the critical phase, with multifractality of the wave function influencing both ground state properties associated with exotic symmetry breaking 18, 24, 25, 26, 27, 28 and the emergence of non-ergodic many-body critical phases 29 at infinite temperature that defies the eigenstate thermalization hypothesis 30, 31, 32, 33.
Figure 1: Schematic for the mechanism of quantum critical states. a-b Density profile of the eigenstates in extended, localized (a) and critical phases (b), which are delocalized in both position and momentum space. Blue circles denote the incommensurately distributed zeros (IDZs), which are key ingredients for the critical states. c Visualization of the self-similar structure characteristic of critical states. d The quantum processor comprising a 2D array of 66 superconducting qubits. Blue spheres represent lattice sites, and blue (orange) bonds mark the nearest-neighbour (long-range) couplings controlled by tunable couplers in the experiment. e Illustration of one-dimensional chain model incorporating long-range couplings. The 2D geometry of superconducting chip facilitates long-range couplings. By activating orange bonds in 2D array and relabeling the system as a 1D chain, nearest-neighbour couplings effectively realize long-range interactions within the redefined 1D configuration.
However, confirming the existence of critical states is significantly challenging. The localized and extended state wave-functions severely fluctuate when the corresponding localization length or correlation length is comparable to the system size. Such fluctuations make localized and extended states resemble critical states. Due to this subtle nature, a precise theoretical framework to characterize such phases has been elusive for decades. Recent breakthroughs, particularly Avila’s global theory, have refined the rigorous characterization 34, 35 of the critical states in quasiperiodic systems 36, 11. This establishes a universal mechanism that critical states can emerge when couplings in one-dimensional (1D) systems are quasiperiodic and feature incommensurately distributed zeros (IDZs) in the thermodynamic limit 11 as illustrated in Fig. 1b, c. Experimentally, identifying critical states presents comparable difficulties. Pioneering experimental efforts have been devoted to probing critical phase 37, 38, 39, 40, 41 in quasiperiodic systems 42, 43, 44, 45, 46, 47, 48, 49, 50. However, the limited system sizes in experiments introduce severe finite-size effects, causing localized and extended states to exhibit behaviors akin to critical states 41. This makes the rigorous experimental confirmation of critical states a challenging task. Despite these obstacles, the elusive observation of transition energies threshold linked to critical states, referred to as anomalous mobility edges (MEs), remains an area ripe for investigation.
In this work, we report the precise characterization of quantum critical states with smoking-gun evidence and further explore their universal behaviors in a mosaic model 11 using a superconducting quantum processor  51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, which reveals universal rigorous mechanism of such critical states as well as the associated mobility edges. With the site-resolved controllability of the coupling strength on each qubit 72, 73, 74, we first observe the delocalized (localized) dynamics and confirm the presence of IDZs in coupling, providing rigorous evidence of the critical (localized) phase for the quasiperiodic mosaic models. We perform the quantum simulation on a superconducting quantum processor comprising a two-dimensional (2D) array of tunably coupled qubits, which enables us to emulate the processes with long-range coupling beyond previous quantum simulations in the one-dimensional chain as illustrated in Fig. 1d. Switching on long-range couplings, we uncover a profound generalized mechanism that the IDZs can protect critical states, as long as they are not completely removed. We further examine the elimination of IDZs in the ergodic dynamics and observe the breakdown of localized-critical phase transition when the IDZs disappear, confirming the IDZs to be a central ingredient of the universal mechanism for stabilizing the non-ergodic critical states. Finally, we observe the anomalous mobility edges separating critical from localized states in the spectra by precisely controlling the coupling coefficients and on-site potentials to the exactly solvable point, which match the results predicted in the thermodynamic limit. This work sets a standard for the unambiguous experimental detection and characterization of quantum critical states, offering a systematic methodology for their observation, and also presents a tunable quantum platform to further explore the novel physics of the critical states.
Figure 2: Characteristic dynamics of localized, critical and extended phases. a-b Dynamics in the localized phase with $\lambda/J=0.25$. c-d Dynamics in the critical phase with $\lambda/J=2.5$. e-f Dynamics in the extended phase with the long-range coupling $J^{L}_{m,n}=\lambda$ and $\lambda/J=2.5$. In a,c,e, the top panels show illustrations for the density profiles of each phases, and bottom panels show the measured dynamics of on-site population $n_{j}(t)$, where the system is initialized in $|\psi(t=0)\rangle=(\prod\sigma^{+}_{6k-5})|0^{\otimes N}\rangle$, with $k$ indexing every sixth site starting from site 1. b,d,f show the measured dynamics of $n_{j}(t)$ with the initial state prepared at the left and right to the zeros in coupling coefficients marked by the stars, highlighting the distinct behaviors of different phases. The parameters of systems are $\lambda/(2\pi)=1~{}\mathrm{MHz}$ (localized phase), $\lambda/(2\pi)=10~{}\mathrm{MHz}$ (critical phase), $\lambda/(2\pi)=J^{L}_{m,n}/(2\pi)=10~{}\mathrm{MHz}$ (extended phase), $J/(2\pi)=4~{}\mathrm{MHz}$, $V_{0}=0$ and $\theta=\pi/5$.
Programmable long-range mosaic model. The model studied in the experiment is effectively described by the Hamiltonian of a 1D chain with $N$ qubits
$H/\hbar$ $=$ $\sum_{j=1}^{N-1}J_{j}(\sigma_{j}^{+}\sigma_{j+1}^{-}+\sigma_{j}^{-}\sigma_{j+1}^{+})+\sum_{j=1}^{N}V_{j}\sigma_{j}^{+}\sigma_{j}^{-}$ (1)
$+$ $\sum_{m,n}J_{m,n}^{L}(\sigma_{m}^{+}\sigma_{n}^{-}+\sigma_{m}^{-}\sigma_{n}^{+}),$
where $\sigma^{+}_{j}$ ($\sigma^{-}_{j}$) represents the raising (lowering) operator for qubits. The first term describes the nearest-neighbour (NN) coupling, with a mosaic quasiperiodic coupling which exhibits IDZs, defined as
$J_{j}=\left\{\begin{array}[]{lll}\lambda&j=1&\bmod~{}2,\\ 2J\cos(2\pi\alpha j+\theta)&j=0&\bmod~{}2.\end{array}\right.$ (2)
Here $\lambda$ and $J$ denote NN coupling strengths, the irrational frequency $\alpha=(\sqrt{5}-1)$/2, and $\theta$ is the phase offset. The second term is a mosaic on-site potential
[…]
with amplitude $V_{0}$. The long-range coupling $J^{L}_{m,n}$ is introduced via tunable couplers $C_{m,n}$ connecting qubits with indices $m$ and $n$ on a 2D lattice configuration, as shown in Fig. 1d, e. This long-range coupling enables the exploration of mechanisms underlying the rigorous critical states and serves as a switch to include extended states in the system. In the limit of short-range coupling $J^{L}_{m,n}=0$ and uniform potential $V_{0}=0$, the system is exactly solvable and exhibits a phase transition between localized and critical phases at $\lambda=J$. The localized phase is obtained for $\lambda<J$, with localization length $\xi=2/\log{|J/\lambda|}$, and the critical phase for $\lambda>J$ (see Supplemental Information). The critical states arise from the generic mechanism of combining the delocalized nature of the states and the IDZs in the coupling coefficients $J_{j}$. Turning on the long-range coupling may remove the IDZs and drive the critical phase into extended phase. On the other hand, introducing a quasiperiodic on-site potential leads to energy-dependent transitions between localized and critical states, where the transition energies $E_{c}$ define the mobility edges. Interestingly, along the high-symmetric line $V_{0}=J$, the system becomes exactly solvable again, allowing for an analytical determination of the localization properties of all eigenstates.
In the experiment, all coupling coefficients and on-site potentials are independently tunable, which we utilize to unravel the mechanism underlying the critical states. We first demonstrate the fundamental organizing principle for the critical states by toggling the long-range coupling in the uniform potential limit $V_{0}=0$, highlighting the key role played by the IDZs for realizing the critical states. Following this, we map out the complete phase diagram for the localized-to-critical phase transition. Finally, we probe the MEs by adjusting the on-site potential, further showcasing the precise tunability of the system.
Figure 3: Localized-critical phase transition and its breakdown. a Measured time-averaged observable $\overline{\mathcal{D}}$ within the time range from 00 to $t_{f}$ for long-ranged coupling $J_{m,n}^{L}=0$, where $\lambda/J=1/4$ ($\lambda/J=2$) are chosen from a deeply localized (critical) phase. b-d Characterization of the localized-critical phase transition for $J_{m,n}^{L}=0$ in (b) and its breakdown in the presence of long-ranged coupling in (c-d), where (c) shows $\overline{\mathcal{D}}$ and time-averaged $\overline{n}_{\mathrm{r}}=\sum_{j>j_{0}}n_{j}$ for different $J_{m,n}^{L}/J$ with $\lambda/(2\pi)=10~{}\mathrm{MHz}$, and (d) gives $\overline{\mathcal{D}}$ and $M(t_{f})$ for different $\lambda/J$ with $J_{m,n}^{L}=2.5J$. The system is initialized in $|\psi(t=0)\rangle=|1\rangle_{14}$, $J/(2\pi)=4~{}\mathrm{MHz}$, and the size of system is $N=24$. The markers represent experimental data, and the solid lines correspond to simulations.
Universal mechanism for critical states. We first identify the key mechanism relating the critical states to the presence of IDZs by studying the characteristic dynamics of different types of states. The time evolution of density population $n_{j}(t)=\lvert\langle\psi(t)|1\rangle_{j}\rvert^{2}$ with $\lvert 1\rangle_{j}=\sigma^{+}_{j}|0^{\otimes N}\rangle$, reveals the unique localization properties of different states as shown in Fig. 2. Here $\lvert\psi(t)\rangle=\exp(-iHt)\lvert\psi(0)\rangle$, and $\lvert\psi(0)\rangle$ is the initial state. Without the on-site modulation and long-range couplings, we observe only the localized and critical dynamics. In particular, for the localized phase with $J>\lambda$, the density profile remains confined as shown in Fig. 2a, b. The localization can be understood as follows. Diagonalizing the dominant $J_{j}=2J\cos(2\pi\alpha j)$ term in Hamiltonian Eq. (1) yields disconnected dimers with quasiperiodically distributed energies, which act as an incommensurate on-site potential. The smaller uniform coupling $J_{j}=\lambda$ connects these dimers, and the system resembles double Aubry-André chain 43 in localization phase. In comparison, for critical phase with $J<\lambda$ in Fig. 2c, d, the quantum states propagate while preserving their local configurations, indicating a critical behavior. Finally, when the long-range coupling is turned on and $J^{L}_{m,n}=\lambda$, the density profile rapidly expands over the entire system (see Fig. 2e, f), erasing the initial configuration and manifesting an extended phase.
The emergence of the critical states is observed uniquely connecting to IDZs, as located in bonds at sites $j=14$ and $j=47$ in our sample, through which the tunneling takes longest time (Fig. 2c,d). The nearly vanishing tunneling leads to the unique uni-side quantum dynamics. The lower panel of Fig. 2 depicts the single-spin evolution with initial states $\lvert\psi(t=0)\rangle=\lvert 1\rangle_{14}$ and $\lvert 1\rangle_{15}$, respectively. For critical phase, we observe that the wave-packet propagates in one side of the bond connecting sites $j=14$ and $15$, with negligible probability in another side (see Fig. 2d). This uni-side propagation quantum dynamics provides a characteristic signature to identify and benchmark IDZs within the system. In contrast, for the long-range coupling $J^{L}_{m,n}=\lambda$, the uni-side quantum dynamics breaks down, with the propagation across the bond at $j=14$ being clearly observed (Fig. 2f). Thus the IDZs disappear and the system turns to extended phase (Fig. 2e). The distinct quantum dynamics associated with the IDZs enables a comprehensive exploration of the entire phase diagram based on the present programmable 2D sample by tuning the long-range couplings.
Figure 4: Mobility edges in the quasiperiodic mosaic lattice. a Measured time-averaged observable $\overline{\mathcal{D}}$, where the system is quenched from the localized state $|\psi_{g}^{+}\rangle$ and zero-energy state $|\psi_{g}^{-}\rangle$ of the Hamiltonian $H$ as $\lambda\to 0$ with $|\psi_{g}^{\pm}\rangle$ $=(|1\rangle_{12}\pm|1\rangle_{13})/\sqrt{2}$. The site $j$ ranging from 1 to $N=24$ and $\theta=\pi/5$. b-c Measured $\overline{\mathcal{D}}$ and $M(t_{f})$ for the system quenched from the superposition of localized and critical state $|\psi^{\phi}_{n}\rangle=(|1\rangle_{12}+e^{i\phi}|1\rangle_{13})/\sqrt{2}$ (b), as well as for states $|\psi^{\pm}_{n}\rangle$ located near the edge (inside the novel mobility edges) and center (outside the novel mobility edges) of the spectrum (c), where $t_{f}=\mathrm{300~{}ns}$ and $|\psi^{\pm}_{n}\rangle=(|1\rangle_{n}\pm|1\rangle_{n+1})/\sqrt{2}$. The dashed lines mark the mobility edges separating the localized and critical states. The markers represent experimental data, and the solid lines correspond to numerical simulations.
[…]
With the long-range couplings, we experimentally discover a novel generalized mechanism that IDZs and critical states are robust to weak $J_{m,n}^{L}$. The transition from critical to extended states occurs only when $J_{m,n}^{L}$ exceeds a threshold, and can be observed by measuring the characteristic uni-side propagation dynamics of critical states. We quantify the observation through the time-averaged density population across IDZ (say at $j=j_{0}$) observed in above section $\overline{n}_{\mathrm{r}}=\sum_{j>j_{0}}n_{j}$. The critical state turns to an extended one when both $\overline{n}_{r}$ and $\overline{\mathcal{D}}$ saturate, indicating the breakdown of uni-side dynamics and the IDZs’ fully disappearing. As shown in Fig. 3c, the small long-range coupling does not ruin the critical states, with $\overline{n}_{r}\approx 0$. Instead, the IDZs of the dominating $J_{j}$-terms are dressed by long-range couplings, and stabilize the critical states in this regime. When the long-range coupling exceeds a threshold $J_{m,n}^{L}/J\gtrsim 0.4$, both $\overline{\mathcal{D}}$ and $\overline{n}_{r}$ saturate as marked by the blue shadow, consistent with the numerical results. In theory, we first model the dressed IDZs due to the next-nearest-neighbor (NNN) couplings $J_{\rm nn}$, and show the critical-extended phase transition analytically by renormalization group method (see Supplementary Material). The transition criterion is given by
$J_{\mathrm{nn}}>\max\big{(}J,\sqrt{J\lambda}\big{)},$ (4)
indicating that critical states are robustly protected by IDZs and critical-extended transition only happens only when NNN couplings exceed a certain threshold. Furthermore, the transition $J_{\rm nn}$ decreases when next-next-nearest-neighbor hopping $\mu$ is involved. When $\mu$ is relatively small, the transition criterion becomes
$J_{\rm nn}>\sqrt{J\max\left(J,\lambda,\mu\right)-\lambda\mu}.$ (5)
For $\lambda=2.5J$, one finds that the transition $J^{c}_{\rm nn}\approx 0.4$ requires to take $\mu\approx 0.93$. While the Eq. (5) is more precise for relatively small $\mu$, this result accounts for the experimental observation that the critical-extended transition occurs at moderate $J_{m,n}^{L}/J$ due to the presence of long-range couplings.
To further illustrate the transition between critical and extended phases, we also measure the phase diagram versus $\lambda$ with fixed long-range coupling $J_{m,n}^{L}$, as shown in Fig. 3d. Turning on a large uniform long-range coupling $J_{m,n}^{L}=2.5J$ always drives the system at any $\lambda$ into an extended phase, while increasing $\lambda$ further delocalizes the system, resulting in larger accumulated $\overline{\mathcal{D}}$ and wider $M(t_{f})$ compared to the case $\lambda=0$. This feature further confirms the transition from critical to extended phase under large $J_{m,n}$. These measurements show the rigorous criteria for quantum critical states: the IDZs in the dominating hopping couplings $J_{j}$ and the delocalized nature in the regime $\lambda>J$. The two criteria guarantee the self-similar and scale-invariance of critical states.
Probing mobility edges. Having established the experimental detection and characterization of critical states, we can probe the new MEs between critical and localized states 11. For this we turn off long-range couplings but switch on the quasiperiodic potential $V_{j}$. From numerical results, we show when $V_{0}/J$ is small, only the edges of the spectrum remain localized, while for $V_{0}\gg J$, most of the spectrum becomes localized (see Supplementary Material). At the high symmetry line $V_{0}=J$, the system is again exactly solvable, yielding MEs at $E_{c}=\pm\lambda$, with critical (localized) states for the region $|E|<|\lambda|$ ($|E|>|\lambda|$). In particular, in this regime the ratio of localized to critical states is $1:1$, which facilitates the probe of MEs between the two types of states.
We measure the MEs for $V_{0}/J=1$ by setting $J/(2\pi)=4$ MHz in the experiment. We start by demonstrating the ME enriched quench dynamics of entangled states near the IDZ. Specifically, we initialize the system in $|\psi_{g}^{\pm}\rangle=(|1\rangle_{12}\pm|1\rangle_{13})/\sqrt{2}$, which are eigenstates of the Hamiltonian for $\lambda\to 0$. We then evolve the initial states under the Hamiltonian at $\lambda/J=1/4$. In Fig. 4a, $\overline{\mathcal{D}}$ of the initially localized state $|\psi^{+}_{g}\rangle$ ($|\psi^{-}_{g}\rangle$) displays a slow (fast) accumulation. Such qualitatively different dynamics of $|\psi_{g}^{\pm}\rangle$ within the same parameter of the system, indicating the presence of ME in the system. We further incorporate more initial states with energies by adjusting the relative phase $\phi$ of the state $\lvert\psi^{\phi}_{g}\rangle=(|1\rangle_{12}+e^{i\phi}|1\rangle_{13})/\sqrt{2}$. As shown in Fig. 4b, $M(t_{f})$ displays a non-trivial pattern of rise-fall as the $\phi$ evolves, and $\overline{\mathcal{D}}$ shows a similar trend. This indicates that the changes in $\phi$ alter the ratio of localized to critical states, leading to distinct dynamics and suggesting the coexistence of these states within the spectrum resulted by the MEs.
Finally, we probe the localized-to-critical transition behavior as the initial energies cross the ME: $|E|=\lambda$ with $\lambda/J=1.5$ and the initial states $\lvert\psi^{\pm}_{n}\rangle=(|1\rangle_{n}\pm|1\rangle_{n+1})/\sqrt{2}$ for various $n$, as shown in Fig. 4c. Here, the initial states are chosen to realize typical energies both inside and outside the MEs, with $n=\{4,8,10,12,16,18\}$ for the plus sign and an additional $n=18$ for the minus sign. When the initial states are prepared as $|E|>\lambda$, the $\overline{\mathcal{D}}$ and $M(t_{f})$ saturates a small number, corresponding to large amount of localized orbitals of such initial states. Conversely, if the initial state is prepared near the center of the spectrum ($|E|<\lambda$), its dynamics displays clear critical characteristics with much larger $M(t_{f})$ accompanied with $\overline{\mathcal{D}}\sim 0.62$. This highlights that $E_{c}=\pm\lambda$ serves as a critical threshold for the energy-dependent spin dynamics spreading, effectively separating the localized and critical orbitals.
Discussion and outlook.
In this work, we experimentally confirmed the presence of quantum critical states in a superconducting quantum simulator with tunable (long-range) couplings and on-site potentials. We discovered in experiment and showed with rigorous theory a novel generalized mechanism of critical states, highlighting the role of incommensurately distributed zeros (IDZs) in stabilizing the critical state. In particular, the critical states survive as long as quasiperiodic zeros are not fully removed by long-range couplings. We further investigate the mobility edges between critical and localized states, revealing enriched spin evolution and energy-dependent quench dynamics near the IDZs.
This study established a standard to rigorously characterize the critical states in experiment, applicable to broad range of quantum simulation and quantum computing systems, and provided a versatile platform to further explore novel quantum physics of critical states, in particular for higher dimensional systems and the many-body regime. For instance, the inter-particle interactions can be introduced by incorporating anharmonicity in the qubits 75, and shall lead to the emergence of novel phases that have not been precisely observed before. The multifractal nature of critical states renders the ground phase of the system highly sensitive to interactions, enriching the exotic symmetry-breaking phases 18, 24, 25, 26, 27, 28. For higher energy states, the non-interacting critical states turn to many-body critical phases 29—a third type of fundamental phase in disordered system, which is different from the thermal phase and many-body localization. Building on the present setting, these highly important quantum many-body phases hold the promise for our next studies.
Beyond the closed system framework, the 2D array of tunably coupled qubits offers high controllability of noise 76, 77, enabling the in-depth study of the robustness of critical states and anomalous MEs against decoherence. Also, by combining our tunable qubits with tailored dissipation, one can explore dissipation-induced critical phases and mobility edges 78, 79. Moreover, the combination of many-body interactions and controlled noise opens up the way to experimentally probe transitions between many-body critical phase in open systems and generic thermalized states. This is an area that remains unexplored both theoretically and experimentally, yet is of fundamental importance.